OpenStudy (anonymous):

One complicated trig identity, Someone help! (tan x) / (sin x- cos x) = (sin^2 x + (sin x)(cos x))/ (cos x - 2 cos ^3 x)

OpenStudy (amistre64):

LHS: sin --- cos ------- sin - cos ---- 1 flip the bottom over and multiply to get...

OpenStudy (amistre64):

sin --------- = RHS cos(sin-cos) RHS: (sin^2 x + (sin x)(cos x))/ (cos x - 2 cos ^3 x) sin(sin + cos) ------------- cos(1 - 2cos^2)

OpenStudy (amistre64):

1 sin + cos tan ------- = tan -------- (sin - cos) 1 - 2cos^2 divide both sides by tan:......

OpenStudy (amistre64):

multiply both side by (-1) to get: -1 sin + cos ------- = --------- sin-cos 2cos^2 -1

OpenStudy (amistre64):

cross multiply....im sure :) 2cos^2 -1 = 2(sin+cos)

OpenStudy (amistre64):

2 cos^2 -sin^2 -cos^2 = 2(sin + cos) cos ^2 - sin^2 = 2(sin+cos)....

OpenStudy (amistre64):

sin+cos(sin - cos) = sin^2 - cos^2 NOT (2)sin+cos

OpenStudy (amistre64):

go back to the cross multiply...

OpenStudy (anonymous):

Thank you for showing me, but unfortunately I'm required to work on one side

OpenStudy (amistre64):

if you can work it out to meet in the middle, then you simply reverse it to get back to the one side.... How you solve it really doesnt matter a hilll of beans :)

OpenStudy (anonymous):

ok, thanks :)

OpenStudy (amistre64):

but let me show you that it DOES work out :) I gots: -(2cos^2 -1) = sin^2 - cos^2 -(2cos^2 -(sin^2 + cos^2)) = sin^2 - cos^2 -(2cos^2 -sin^2 -cos^2) = sin^2 - cos^2 -(cos^2 - sin^2) = sin^2 - cos^2 sin^2 - cos^2 = sin^2 - cos^2 (proofed)

OpenStudy (anonymous):

I'm kinda confused now...

OpenStudy (amistre64):

who requires you ta work only one side?

OpenStudy (anonymous):

my professor requires us to work one side to equal other side

OpenStudy (amistre64):

since both sides are equally "messed up"; it would be best to "simplify" both sides till they are equal. I you can only work with one die, you are pretty much tieing your hands behind your back.....

OpenStudy (amistre64):

my keyboard hates me :)

OpenStudy (anonymous):

ok, its fine lol.. i appreciate the help, it would be easier to work both sides.. wish i could